OpenStudy (anonymous):

Kaley bought an oven from a rental center for $728.44. She makes 12 monthly payments of $74.91 with her credit card. The rental center charges $2.75 for every purchase made with a credit card. She also had to pay late fees in the amount of $35 four different times. What is the total cost of the oven? $1,071.92 $931.92 $898.92 $1,038.92

OpenStudy (australopithecus):

this is simple addition read it out and write down all the money she spent, then take the sum

OpenStudy (australopithecus):

If you have any questions please ask

OpenStudy (anonymous):

= 12(74.91) + 12(2.75) + 4(35) =898.92 + 33 + 140 = 1071.92 answer is A

OpenStudy (australopithecus):

@Omistone, this isn't an answer site. You are suppose to help people arrive at the answer themselves, not give it to them.

OpenStudy (anonymous):

I was just stepping the person through the process, now if they ever needs to figure out a similar question, they will know how to go about it. Most people I help thank me for the way I explain the question logically and simply to help them come to an understanding. I understand where you are coming from, @Australopithecus , but I try to equip people with the skills to tackle these questions for themselves. If they ever have any questions or do not understand how I jump from one conclusion to the next, they are free to ask and I will explain it until they have a clear understanding of the concepts involved. Thanks for the heads-up though.